OpenStudy (anonymous):

What are the x- and y- intercepts of the following equation: 8x - 4y = 12 A. (0, 3), (-3/2, 0) B. (0, -3), (3/2, 0) c.(-3, 0), (0, 3/2) d.(3, 0), (3/2, 0)

OpenStudy (anonymous):

8x-0=12 8x=12 x = 3/2

OpenStudy (anonymous):

0-4y=12 -4y=12 y=-3

OpenStudy (anonymous):

c.(-3, 0), (0, 3/2) <-- will be your answer !

OpenStudy (anonymous):

:D